Three blocks on a frictionless surface are connected by massless strings, with M1 = 1.50 kg, M2 = 2.60 kg, and M3 = 3.40 kg. Due to the force F acting on M3, as shown, the system accelerates to the right. Given that T1 is 3.10 N, calculate T2
Answer #1

assuming that the system is something like this

M1-----T1------M2-----T2------M3----F

since the acceleration of all the blocks are the same, let it be a.

T1=M1*a

or a=3.1/1.5

=2.0667

for the block 2,

T2-T1 = M2*a

or T2 - 3.1 = 2.6*2.0667

or T2=8.47 N